# Convergence of Unadjusted Langevin in High Dimensions: Delocalization of Bias

### v1

@@ 2.3. Examples of sparse graphical models
It is not immediately apparent that given constant values of $\alpha$ and $\beta$ such that $\alpha\leq\beta$, there exists a potential function $V$ for any dimension $d\in\mathbb{N}$ with $\alpha I\preceq\nabla^{2}V(x)\preceq\beta I$ at any point $x\in\mathbb{R}^{d}$, and where $s_{k}$ grows at most polynomially in $k$, as required in Theorem 2.4. As the following example shows, a natural class of examples of such potentials arises from Laplacians of sparse graphs.
@@
This matrix can be seen as $-\Delta+\lambda(x)I$, where $\Delta$ is the discretization of the 1D Laplace operator with homogeneous Dirichlet boundary conditions. This operator satisfies $0\preceq-\Delta\preceq 4I$. Therefore, if $V$ is a potential such that $\nabla^{2}V(x)$ is equal to the above matrix and $\lambda$ is a bounded function satsifying $\min_{x\in\mathbb{R}^{d}}\lambda(x)=\alpha>0$, then we have $s_{k}=\max\{2k+1,d\}$ and $\alpha I\preceq\nabla^{2}V(x)\preceq\beta I$ with $\beta=4+\max_{x\in\mathbb{R}^{d}}\lambda(x)$, in any dimension $d$. Thus this model satisfies the assumption in Theorem 2.4 with $n=1$.
@@ 3. Discussions and Generalization with Asymptotic Arguments
Below we calculate the first order formula for the bias of an observable $f$. We assume that $f$ is sufficiently regular so that the Taylor expansions and integration by parts used in the Proof of Proposition 3.1 below are valid. In this section, we use the notation $A=o(h)$ to mean that the term $A$ is a higher order term than $h$ as $h\to 0$.
@@
Without loss of generality, assume $\int f\pi=0$. Then it holds that
@@
The proof of the formula can be found in Appendix D. Let us make some observations regarding Proposition 3.1. If $\pi$ is Gaussian, then $\int f(\Delta\log\pi)\pi=0$ since $\int f\pi=0$. We have
@@
In fact, the above argument can be further generalized to $\pi$ that is a perturbation of Gaussians; see the following Proposition 3.2, and its proof is in Appendix D.
@@
Let $\pi\propto\exp(-V(x))\propto\mathcal{N}(x;m,\Sigma)\exp(-U(x))$ where $\mathcal{N}(x;m,\Sigma)$ is the density of a Gaussian with mean $m$ and covariance $\Sigma$. Assume $\pi$ is centered, and $\int(\Delta U)^{2}\pi\leq C_{0}$ for some constant $C_{0}$ independent of $d$. Suppose $f$ depends only on $K$ coordinates of $x\in\mathbb{R}^{d}$, then the first-order term of the bias depends on $K$, not $d$. In particular, for the observable $f(x)=\sum_{k=1}^{K}x^{(k)}$, under the additional assumption that $\pi$ satisfies a Poincaré inequality with a positive dimension-independent constant, we have
@@ 4. Conclusions
Our discussion hinges on a delocalization effect of the bias: the overall bias of the algorithm is delocalized to a comparable order across each one-dimensional marginal distribution. Thus, to obtain accurate low-dimensional marginals, one does not suffer from a poor dependence on the full dimension $d$. The delocalization effect is rigorously shown in this paper for Gaussian distributions and distributions with certain sparsity structures. Our error bounds have a favorable dependence on $d$; however, the dependence on the condition number $\beta/\alpha$ may potentially be improved. We also provide counterexamples of a rotated product measure for which the bias is not delocalized. Extending the delocalization analysis to other broader classes of distributions is of future interest. In fact, our asymptotic arguments based on Poisson equations provided further insights into how the first-order terms of the bias of general observables scale with dimension.
More broadly, it is an important question to understand how the complexity of other unadjusted MCMC algorithms scales with dimension when the quantities of interest depend solely on low-dimensional marginals. Addressing this question can provide valuable insights to help practitioners select appropriate algorithms and understand their computational complexity when dealing with high-dimensional sampling problems.
Acknowledgments We thank Nawaf Bou-Rabee and Aaron Dinner for helpful discussions. Y. Chen is supported in part by the Office of Naval Research project under award N00014-22-1-2728 (PI: Benjamin Peherstorfer) and Vannevar Bush award “Mathematical Foundations and Scientific Applications of Machine Learning” (PI: Eric Vanden-Eijnden). J. Weare and X. Cheng are supported in part by National Science Foundation award DMS-2054306. J. Niles-Weed is supported in part by National Science Foundation awards DMS-2210583 and DMS-2339829.
@@ Appendix A Proofs for Log-Concave Product Measures
(A.1) $W_{2,\ell^{\infty}}(\rho_{(k+1)h},\pi)\leq qW_{2,\ell^{\infty}}(\rho_{kh},\pi)+\beta\sqrt{(\frac{8\beta^{2}}{3\alpha}h^{4}+8h^{3})\log(2d)}\,.$
@@
$W_{2,\ell^{\infty}}(\pi_{h},\pi)\leq\frac{\beta}{\alpha}\sqrt{(\frac{8\beta^{2}}{3\alpha}h^{2}+8h)\log(2d)}\,.$
Furthermore, if $h\leq\frac{\alpha}{\beta^{2}}$, we get
@@
$\leq$ $\int_{kh}^{(k+1)h}\left(2\mathbb{E}[|\int_{kh}^{t}\nabla V(Y_{s}){\rm d}s|_{\infty}^{2}]+2\mathbb{E}[\sqrt{2}B_{t-kh}|_{\infty}^{2}]\right){\rm d}t$
$\leq$ $\int_{kh}^{(k+1)h}2(t-kh)\int_{kh}^{t}\mathbb{E}[|\nabla V(Y_{s})|_{\infty}^{2}]{\rm d}s{\rm d}t+\int_{kh}^{(k+1)h}2\mathbb{E}[\sqrt{2}B_{t-kh}|_{\infty}^{2}]{\rm d}t$
$\leq$ $2\int_{kh}^{(k+1)h}(t-kh)^{2}\mathbb{E}_{\pi}[|\nabla V(Y)|_{\infty}^{2}]{\rm d}t+\int_{kh}^{(k+1)h}16(t-kh)\log(2d){\rm d}t$
@@
Then by Proposition 2.3, we have $\sqrt{\mathbb{E}_{\pi}[|\nabla V|_{\infty}^{2}]}\leq 2\frac{\beta}{\sqrt{\alpha}}\sqrt{\log(2d)}$. Combining all the bounds above, we get
$\sqrt{\mathbb{E}[|X_{(k+1)h}-Y_{(k+1)h}|_{\infty}^{2}]}\leq q\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}+\frac{\beta}{\alpha}\sqrt{(\frac{8\beta^{2}}{3\alpha}h^{4}+8h^{3})\log(2d)}\,.$
@@
$W_{2,\ell^{\infty}}(\rho_{(k+1)h},\pi)\leq qW_{2,\ell^{\infty}}(\rho_{kh},\pi)+\frac{\beta}{\alpha}\sqrt{(\frac{8\beta^{2}}{3\alpha}h^{2}+8h)\log(2d)}\,.$
Iterating this inequality leads to the bound on the $W_{2,\infty}$ bias. Furthermore under the assumption $h\leq\frac{\alpha}{\beta^{2}}$, we get
$W_{2,\ell^{\infty}}(\pi_{h},\pi)\leq\frac{\beta}{\alpha}\sqrt{(\frac{8}{3}+8)h\log(2d)}\leq\frac{4\beta}{\alpha}\sqrt{h\log(2d)}\,.$
The proof is complete. ∎
@@ Appendix B Proofs for Gaussian Distributions
Setting $Y=(\Sigma^{1/2}-\Sigma^{1/2}_{\infty})Z,$ we see that $Y\sim\mathcal{N}(0,\sigma_{i}^{2})$, where $\sigma_{i}^{2}=((\Sigma^{1/2}-\Sigma^{1/2}_{\infty})^{2})_{ii}\leq|\Sigma^{1/2}-\Sigma^{1/2}_{\infty}|_{2}^{2}$. In particular, the entries of $Y$ are $|\Sigma^{1/2}-\Sigma^{1/2}_{\infty}|_{2}^{2}$-subgaussian.
@@ C.1. Sparsity of the propagator of unadjusted Langevin
    (C.5) $|J_{N}|_{\infty}\leq\beta(\sqrt{s_{r}}q^{N-1}+\sqrt{d}\exp(-r))$
@@
is nonzero if and only if $j\in\mathsf{N}_{r}(i)$. Since $\max_{1\leq i\leq d}|\mathsf{N}_{r}(i)|=s_{r}$, we will have at most $s_{r}$ nonzero entries in each row. This implies that the matrix is $s_{r}$-sparse.
@@
$\leq\sqrt{s_{r}}\sqrt{\sum_{j=1}^{N}|(P_{N})_{ij}|^{2}}+\sqrt{d}\exp(-r)$
@@
Proof for (iii). Note that $J_{N}=\left(\int\nabla^{2}V\,{\rm d}\nu_{0}\right)H_{N}$. Following the proof for (ii), we have that when $r\geq e^{2}Nh\beta$,
(C.12) $\left|\left(\int\nabla^{2}V\,{\rm d}\nu_{0}\right)\sum_{k=r+1}^{N}(-1)^{k}h^{k}\sum_{1\leq i_{1}<...<i_{k}\leq N}A_{i_{1}}...A_{i_{k}}\right|_{2}$
$\leq$ $\beta\sum_{k=r+1}^{N}(\frac{eNh\beta}{k})^{k}\leq\beta\sum_{k=r+1}^{N}\exp(-k)\leq\beta\exp(-r)\,.$
@@
(C.13) $\sum_{j=1}^{N}|(J_{N})_{ij}|\leq\beta\sqrt{s_{r}}q^{N}+\sqrt{d}\beta\exp(-r)\,.$
Therefore, $|J_{N}|_{\infty}\leq\beta(\sqrt{s_{r}}q^{N-1}+\sqrt{d}\exp(-r))$ for $r\geq e^{2}Nh\beta$. Taking $r_{N}=\lceil e^{2}Nh\beta+\log\sqrt{d}\rceil$, we get $|J_{N}|_{\infty}\leq 2\beta\sqrt{s_{r_{N}}}q^{N}$. The proof is complete. ∎
@@ C.2. A multistep coupling argument
where we applied (iii) of Proposition C.1 to the the matrix $H_{k}\nabla^{2}V(uY_{t}+(1-u)Y_{(k-1)h})$. As a summary, we get
@@ C.3. Convergence bounds
By the Bakry-Émery criterion, for distribution $\pi$ that is $\alpha$-strongly log concave, $\pi$ satisfies the log-Sobolev inequality with constant $1/\alpha$. Then, by Herbst’s argument [42], we have
(C.26) $\mathbb{E}_{\pi}[\exp(\lambda(Y^{(i)}-\mathbb{E}_{\pi}Y^{(i)}))]\leq\exp(\frac{\lambda^{2}}{2\alpha})\,.$
Furthermore, since $\nabla V$ is $\beta$-Lipschitz, Herbst’s argument leads to
(C.27) $\mathbb{E}_{\pi}[\exp(\lambda(\nabla_{i}V(Y)-\mathbb{E}_{\pi}\nabla_{i}V(Y))]\leq\exp(\frac{\lambda^{2}\beta^{2}}{2\alpha})\,,$
where $\nabla_{i}V$ is the $i$-th component of the vector $\nabla V$. This implies that the distribution of $\nabla_{i}V(Y)$ has a sub-Gaussian tail. Thus, by Lemma B.1, we get
(C.28) $\sqrt{\mathbb{E}_{\pi}[|\nabla V(Y)-\mathbb{E}_{\pi}[\nabla V(Y)]|_{\infty}^{2}]}\leq 2\frac{\beta}{\sqrt{\alpha}}\sqrt{\log(2d)}\,.$
@@
Proof of Proposition 2.4.
We will use Theorem 2.2. Recall the definition $r_{N}=\lceil e^{2}Nh\beta+\log\sqrt{d}\rceil$ and $q=\exp(-h\alpha)$. There is a universal upper bound on $s_{i}$, given by $s_{i}\leq d$. Choose $N=\lceil\frac{\log(4\sqrt{d})}{h\alpha}\rceil$, which leads to $22q^{N}\sqrt{s_{r_{N}}}\leq 1/$ 2. Based on Theorem 2.2, it remains to calculate the bound on $\sum_{i=1}^{N}q^{i-1}\sqrt{s_{r_{i}}}$.
@@
(C.33) $h^{2}\sqrt{\mathbb{E}_{\pi}[|\nabla V(Y)|_{\infty}^{2}]}+3h^{3/2}\sqrt{\log(2d)}\lesssim\left(\frac{\beta}{\sqrt{\alpha}}h^{2}+h^{3/2}\right)\sqrt{\log(2d)}\,.$
Using the fact that $h\leq\frac{\alpha}{\beta^{2}}$ and combining the above two inequalities leads to the final result. ∎

### v2

@@ 2.3. Examples of sparse graphical models
We show that there exist non-trivial examples satisfying our assumptions. Specifically, for given constants $\alpha$ and $\beta$ with $\alpha\leq\beta$, there exists a potential function $V$ defined for any dimension $d\in\mathbb{N}$ such that $\alpha I\preceq\nabla^{2}V(x)\preceq\beta I$ for all $x\in\mathbb{R}^{d}$, and $s_{k}$ grows at most polynomially in $k$, as required in Theorem 2.4. A natural class of such potentials arises from Laplacians of sparse graphs, as illustrated in the following example.
@@
This matrix can be seen as $\Delta+\lambda(x)I$, where $\Delta$ is the discretization of the 1D Laplace operator with homogeneous Dirichlet boundary conditions. This operator satisfies $0\preceq-\Delta\preceq 4I$. Therefore, if $V$ is a potential such that $\nabla^{2}V(x)$ is equal to the above matrix and $\lambda$ is a bounded function satisfying $\min_{x\in\mathbb{R}^{d}}\lambda(x)=\alpha>0$, then we have $s_{k}=\max\{2k+1,d\}$ and $\alpha I\preceq\nabla^{2}V(x)\preceq\beta I$ with $\beta=4+\max_{x\in\mathbb{R}^{d}}\lambda(x)$, in any dimension $d$. Thus this model satisfies the assumption in Theorem 2.4 with $n=1$.
@@ 3. Discussions and Generalization with Asymptotic Arguments
Below we calculate the first order formula for the bias of an observable $f$. Here, we use the notation $A=o(h)$ to mean that the term $A$ is a higher order term than $h$ as $h\to 0$.
@@
Without loss of generality, assume $\int f\pi=0$. Then, under regularity assumptions provided in Appendix D, it holds that
@@
The proof of the formula can be found in Appendix D.
Let us make some observations regarding Proposition 3.1. If $\pi$ is Gaussian, then $\int f(\Delta\log\pi)\pi=0$ since $\int f\pi=0$. We have
@@
In fact, the above argument can be further generalized to $\pi$ that is a perturbation of Gaussians; see the following Proposition 3.2, and its proof in Appendix D.
@@
Let $\pi\propto\exp(-V(x))\propto\mathcal{N}(x;m,\Sigma)\exp(-U(x))$ where $\mathcal{N}(x;m,\Sigma)$ is the density of a Gaussian with mean $m$ and covariance $\Sigma$. Assume $\pi$ is centered, and $\int(\Delta U)^{2}\pi-(\int\Delta U\pi)^{2}\leq C_{0}$ for some constant $C_{0}$ independent of $d$. Suppose $f$ depends only on $K$ coordinates of $x\in\mathbb{R}^{d}$, then the first-order term of the bias depends on $K$, not $d$. In particular, for the observable $f(x)=\sum_{k=1}^{K}x^{(k)}$, under the additional assumption that $\pi$ satisfies a Poincaré inequality with a positive dimension-independent constant, we have
@@ 4. Conclusions
The delocalization of bias effect is rigorously shown in this paper for Gaussian distributions and distributions with certain sparsity structures. Our error bounds have a favorable dependence on $d$; however, the dependence on the condition number $\beta/\alpha$ may potentially be improved. We also provide counterexamples of a rotated product measure for which the bias is not delocalized.
Our work represents a first step in understanding the delocalization of bias phenomenon. Though our sparsity assumption is inspired by physical systems such as statistical physics models and molecular dynamics simulations, it does not fully address those problems. For example, extensions to systems with decaying but long-range interactions and non-logconcave densities should be considered. In fact, our asymptotic arguments based on Poisson equations provide further insights into how the first-order terms of the bias of general observables scale with dimension.
More broadly, it is important to understand how the complexity of other unadjusted MCMC algorithms scales with dimension when the quantities of interest depend solely on low-dimensional marginals. Addressing this question can provide valuable insights to help practitioners select appropriate algorithms and understand their computational complexity when dealing with high-dimensional sampling problems.
Acknowledgments We thank Nawaf Bou-Rabee and Aaron Dinner for helpful discussions. We thank Sinho Chewi for pointing us to [1, Theorem 1.2]. Y. Chen is supported in part by the Office of Naval Research project under award N00014-22-1-2728 (PI: Benjamin Peherstorfer) and Vannevar Bush award “Mathematical Foundations and Scientific Applications of Machine Learning” (PI: Eric Vanden-Eijnden). J. Weare and X. Cheng are supported in part by National Science Foundation awards DMS-2054306 and DMS-2425899. J. Niles-Weed is supported in part by National Science Foundation awards DMS-2210583 and DMS-2339829.
@@ Appendix A Proofs for Log-Concave Product Measures
(A.1) $W_{2,\ell^{\infty}}(\rho_{(k+1)h},\pi)\leq qW_{2,\ell^{\infty}}(\rho_{kh},\pi)+\beta\sqrt{(\frac{8\beta}{3}h^{4}+8h^{3})\log(2d)}\,.$
@@
$\leq$ $\int_{kh}^{(k+1)h}\left(2\mathbb{E}[|\int_{kh}^{t}\nabla V(Y_{s}){\rm d}s|_{\infty}^{2}]+2\mathbb{E}[|\sqrt{2}B_{t-kh}|_{\infty}^{2}]\right){\rm d}t$
$\leq$ $\int_{kh}^{(k+1)h}2(t-kh)\int_{kh}^{t}\mathbb{E}[|\nabla V(Y_{s})|_{\infty}^{2}]{\rm d}s{\rm d}t+\int_{kh}^{(k+1)h}2\mathbb{E}[|\sqrt{2}B_{t-kh}|_{\infty}^{2}]{\rm d}t$
[…]
@@
Then by Proposition 2.3, we have $\sqrt{\mathbb{E}_{\pi}[|\nabla V|_{\infty}^{2}]}\leq 2\sqrt{\beta\log(2d)}$. Combining all the bounds above, we get
$\sqrt{\mathbb{E}[|X_{(k+1)h}-Y_{(k+1)h}|_{\infty}^{2}]}\leq q\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}+\beta\sqrt{(\frac{8\beta}{3}h^{4}+8h^{3})\log(2d)}\,.$
@@
$W_{2,\ell^{\infty}}(\rho_{(k+1)h},\pi)\leq qW_{2,\ell^{\infty}}(\rho_{kh},\pi)+\beta\sqrt{(\frac{8\beta}{3}h^{4}+8h^{3})\log(2d)}\,.$
Iterating this inequality leads to the bound on the $W_{2,\ell^{\infty}}$ bias. Furthermore under the assumption $h\leq\frac{1}{\beta}$, we get
$W_{2,\ell^{\infty}}(\pi_{h},\pi)\leq\frac{\beta}{\alpha}\sqrt{(\frac{8\beta h^{2}}{3}+8h)\log(2d)}\leq\frac{4\beta}{\alpha}\sqrt{h\log(2d)}\,.$
The proof is complete.
@@ Appendix B Proofs for Gaussian Distributions
Setting $Y^{\prime}=(\Sigma^{1/2}-\Sigma^{1/2}_{\infty})Z,$ we see that $Y^{\prime}_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$, where $\sigma_{i}^{2}=((\Sigma^{1/2}-\Sigma^{1/2}_{\infty})^{2})_{ii}\leq|\Sigma^{1/2}-\Sigma^{1/2}_{\infty}|_{2}^{2}$. In particular, the entries of $Y^{\prime}$ are $|\Sigma^{1/2}-\Sigma^{1/2}_{\infty}|_{2}^{2}$-subgaussian.
@@ C.1. Sparsity of the propagator of unadjusted Langevin
    (C.5) $|J_{N}|_{\infty}\leq\beta(\sqrt{s_{r}}q^{N}+\sqrt{d}\exp(-r))$
@@
is nonzero if $j\in\mathsf{N}_{r}(i)$, or if there exists an $k$ such that $i\sim\mathsf{N}_{r}(k),j\sim\mathsf{N}_{r}(k)$. Since $\max_{1\leq i\leq d}|\mathsf{N}_{2r}(i)|=s_{r}$, we will have at most $s_{r}$ nonzero entries in each row. This implies that the matrix is $s_{r}$-sparse.
@@
$\leq\sqrt{s_{r}}\sqrt{\sum_{j=1}^{d}|(P_{N})_{ij}|^{2}}+\sqrt{d}\exp(-r)$
@@
Proof for (iii). Note that $J_{N}=\left(\int\nabla^{2}V\,{\rm d}\nu_{0}\right)P_{N}$. Following the proof for (ii), we have that when $r\geq e^{2}Nh\beta$,
(C.12) $\left|\left(\int\nabla^{2}V\,{\rm d}\nu_{0}\right)\sum_{k=r+2}^{N}(-1)^{k}h^{k}\sum_{1\leq i_{1}<...<i_{k}\leq N}A_{i_{1}}...A_{i_{k}}\right|_{2}$
$\leq$ $\beta\sum_{k=r+2}^{N}(\frac{eNh\beta}{k})^{k}\leq\beta\sum_{k=r+2}^{N}\exp(-k)\leq\beta\exp(-r)\,.$
@@
(C.13) $\sum_{j=1}^{d}|(J_{N})_{ij}|\leq\beta\sqrt{s_{r}}q^{N}+\sqrt{d}\beta\exp(-r)\,.$
Therefore, $|J_{N}|_{\infty}\leq\beta(\sqrt{s_{r}}q^{N}+\sqrt{d}\exp(-r))$ for $r\geq e^{2}Nh\beta$. Taking $r_{N}=\lceil e^{2}Nh\beta+\log\sqrt{d}\rceil$, we get $|J_{N}|_{\infty}\leq 2\beta\sqrt{s_{r_{N}}}q^{N}$. The proof is complete. ∎
@@ C.2. A multistep coupling argument
where we applied (iii) of Proposition C.1 to the the matrix $H_{k}\nabla^{2}V(uY_{t}+(1-u)Y_{(k-1)h})$. Indeed, the bound will depend on $s_{r_{1}}$ if we use the argument in Proposition C.1 directly. Here, we overestimate the bound by bounding $s_{r_{1}}$ by $s_{r_{2}}$, which helps organize the terms in a unified way.
As a summary, we get
@@ C.3. Convergence bounds
Under the assumption, it holds that $\nabla V$ is $\sqrt{\beta}$ sub-Gaussian under $\pi$ [1, Theorem 1.2]. Thus, by Lemma B.1, we get
(C.26) $\sqrt{\mathbb{E}_{\pi}[|\nabla V(Y)-\mathbb{E}_{\pi}[\nabla V(Y)]|_{\infty}^{2}]}\leq 2\sqrt{\beta\log(2d)}\,.$
@@
Proof of Theorem 2.4.
We will use Theorem 2.2. Recall the definition $r_{N}=\lceil e^{2}Nh\beta+\log\sqrt{d}\rceil$ and $q=\exp(-h\alpha)$. There is a universal upper bound on $s_{i}$, given by $s_{i}\leq d$. Choose $N=\lceil\frac{\log(4\sqrt{d})}{h\alpha}\rceil$, which leads to $2q^{N}\sqrt{s_{r_{N}}}\leq 1/2$. Based on Theorem 2.2, it remains to calculate the bound on $\sum_{i=1}^{N}q^{i-1}\sqrt{s_{r_{i}}}$.
@@
(C.31) $h^{2}\sqrt{\mathbb{E}_{\pi}[|\nabla V(Y)|_{\infty}^{2}]}+3h^{3/2}\sqrt{\log(2d)}\lesssim\left(\sqrt{\beta}h^{2}+h^{3/2}\right)\sqrt{\log(2d)}\,.$
Using the fact that $h\leq\frac{1}{\beta}$ and combining the above two inequalities leads to the final result. ∎